A 2,560 mL aqueous sample contained 5.96 g of sodium hydroxide, NaOH.

What is the pH of the solution at 25oC.      

(Na = 22.98 g/mol; O=16.0 g/mol, H=1.00 g/mol)

Conversion

1000.0 mL = 1 L

Molarity = moles/Liter
Answer #1

Molarity of sodium hydroxide = number of moles / volme of solution inliter

number of mole = weight in gram / molecular wight

weight in gram of NaOH = 5.96 g

Molecular weight of NaOH = 40

number of mole = 5.96 / 40 = 0.149 mole

molarity = 0.149 / 2.560 = 0.0582 mole/L

pOH = -log(OH-) = -log (0.0582) = 1.24

pH = 14- P(OH) = 14 - 1.24 = 12.76

Thus pH of this solution is 12.76 at 25oC.